OpenStudy (pulsified333):

The linear function f with values f(−6)=8 and f(5)=3 is f(x)=

OpenStudy (pulsified333):

@ganeshie8

imqwerty (imqwerty):

a linear function is of this form-> \(f(x) =ax+b\) where \(a\) and \(b\) are constants. \(f(-6)=a(-6)+b\) and u know that \(f(-6)=8\) so \(8=-6a+b\) <--------------equation 1 and \(f(5)=a(5)+b\) and u knw that \(f(5)=3\) so \(3=5a+b\) <---------------equation 2 now you got 2 equations solve then to get \(a\) and \(b\) then put the values of \(a\) and \(b\) in this-> \(f(x)=ax+b\) this is the required function